Find the angles of a triangle if one angle exceeds a second angle by 45 degrees and is three times as large as the third angle.
A woman is 5 times as old as her youngest daughter. How old is each now if the sum of their ages is 50?
A motor boat travels at the rate of 50kph down the river. If the current of the river is 5kph, how far can it travel up the river in 2 hours?
The age of a man is 10 years more than three times his brother's age. Find the age of each if 5 years from now he will be 4 years more than twice his brother's age.
A reservoir can be filled by one pipe in 6 hrs and by another pipe in 8 hrs. It can be emptied by a third pipe in 10 hrs. Starting empty, how long will it take to fill the reservoir if all pipes are open?

Answer by mathstutor494(120)

Let one angle be x
So second angle = x-45
and third angle = x/3
As sum of all the 3 angles add up to 180 deg
Therefore x+(x-45)+(x/3) = 180
> x = 96.42
Ans : Angle of the triangles are 96.42, 51.42 and 32.14